OpenStudy (anonymous):

Increasing the average speed of a car by 12 mph results in a 189 mile trip taking an hour less than before. What was the original average speed of the car?

OpenStudy (anonymous):

46 I think

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

189/x=189/(x+12)+1 189(x+12)=189(x)+x2+12x 189x+2268=189x+x2+12x x2+12x-2268=0 (x+58)(x-46)=0 x=46 or -58 Throwing out the negative answer, we get the original speed of the car to be 46 mph..